OpenStudy (anonymous):

cos10-sin10/cos10+sin10

OpenStudy (gorv):

\[\frac{ \cos 10- \sin10 }{ \cos10+\sin10 }*\frac{ \cos10-\sin10 }{ \cos10-\sin10 }\]

OpenStudy (gorv):

\[\frac{ (\cos10-\sin10)^2 }{ (co^210-\sin^210) }\]

OpenStudy (gorv):

\[\frac{ \cos^210+\sin^210-2*\cos10*\sin10 }{ \cos(2*10) }\]

OpenStudy (gorv):

2cosxsinx=sin2*x cos^2x-sin^2x=cos2*x sin^2x+cos^2x=1

OpenStudy (gorv):

use these and try this now

OpenStudy (anonymous):

1

OpenStudy (gorv):

1 ???

OpenStudy (gorv):

dont think ...just solve

OpenStudy (anonymous):

hlo gorv may i knw wher r u frm